### v1

@@ 1 Introduction
In this section, we review the main changes introduced by 5G networks, focusing on the network core, including the network functions for both the control plane and the user plane, as well as the isolation mechanisms and innovations that differentiate it from previous generations. We also present the main open-source 5G applications. The goal is to understand the operation of 5G, with an emphasis on network isolation, roviding the foundation for building an adequate 5G network for experimentation. Finally, we present related work on the topic of 5G network slice isolation.
@@ 2.6.1 UERANSIM
UERANSIM is a simulator for user equipment (UE) and radio access network (RAN) in 5G networks, designed as an open-source solution that simulates both 5G mobile devices and base stations. The project aims to enable testing and research on 5G networks. However, UERANSIM does not offer a complete implementation of the 5G physical layer. The 5G-NR radio interface is only partially implemented and simulated using the UDP protocol [18]. The simulator includes the components nr-gnb and nr-ue, which represent, respectively, the gNodeB (gNB) and the UE. When establishing a connection with the 5G network, the UE creates a network interface responsible for tunneling with the user plane function (UPF).
@@ 2.7 Tool Comparisons
2.8 Related Work
The study described in [20] compares various test environments that utilize 5G slicing, establishing two categories of criteria. The primary criteria require the environment to support technologies such as SDN, NFV, and cloud computing, ensuring flexibility and dynamism. It is crucial that the testbed provides dynamic management, orchestration, and monitoring capabilities (MANO). The environment must span multiple network domains to enable end-to-end (E2E) slicing, even if only partially. Support for multi-tenancy is essential, allowing a single infrastructure to serve multiple tenants, reflecting the capabilities designed for 5G networks.
In contrast, the secondary criteria introduce additional features for slicing testbeds, such as support for multiple radio access technologies, E2E network slicing, multi-location support, machine learning integration, and open-source characteristics. These features enable the implementation of realistic scenarios, resource optimization, and the promotion of innovation through accessible and collaborative platforms [20].
The POSENS protocol, as described in [21], offers an open-source solution for end-to-end network slicing, including both RAN and core. This protocol adopts a shared RAN approach that recognizes different slices, making modifications to the upper layers of network protocol stacks, specifically PDCP and RRC, in both eNB and UE. Using a load balancer (LB) node, packets are analyzed during tunneling between the radio and the user plane, allowing them to be redirected according to specific rules to a new interface and thus forwarded to the appropriate UPF. At this forwarding point, bandwidth limitations are applied to enable traffic prioritization across slices [22].
As detailed in [9], a network slicing test environment was developed using slice selection through the single network slice selection assistance information (NSSAI). To ensure network slice isolation, containers are used, each hosting a specific network core. The environment’s evaluation focused exclusively on the connection point between normal LTE UEs and those with NSSAI capabilities. The aspect of network slice isolation was not addressed in the evaluations conducted.
According to the study described in [23], an approach was developed in which the 5G core is virtualized and containerized, implemented in a cloud infrastructure. This configuration uses a Cisco router and separates into two VLANs to enable distinct connections between the eNB and the user plane. Success in using the cloud infrastructure was reported, and as a next step, the study proposes positioning both the radio and the core at the cloud edge to improve performance and reduce latency.
Concerning slice isolation, [20] observes that the evaluated testbeds do not ensure efficient isolation, offering only partial isolation or limiting slicing to a specific protocol layer. Research such as that of [24] uses Wireguard, a secure network tunnel for Linux that operates at layer 3 and is designed to be a more secure, faster, and easier-to-use alternative compared to IPsec and OpenVPN. Although this work ensures network slice security, it does not guarantee traffic isolation between slices. Despite the various studies analyzed, it is concluded that efficient isolation remains a challenge and a rich area for future research.
@@ 4.1 Experimentation Methodology
During the experiments, several UEs (or pods) are executed over a 20-minute period. Initially, to represent the high-priority slice (Slice 1), 4 pods are started simulating 4 cameras streaming high-quality video at an operating room. Video is streamed over the edge-located UPF and delivered to a simulated cloud service (iPerf server). UEs connected to lower-priority slices (Slices 2 to 5) serve as background traffic and generate the same traffic patterns as the priority ones. In Slices 2 to 5, at every 4 minutes of experimentation, a new UE is added, while the pods are allocated to the node corresponding to the slice, as illustrated in Figure 3. Video transmission is simulated using the isochronous [^6] https://manpages.debian.org/bullseye/iperf/iperf.1.en.html of IPerf, configured with 60 frames per second at a bitrate of 45 Mbits per second, and no variation in frame size.
@@ 4.8 Experiments with CPU limitation, bandwidth limitation, and prioritization
Experiments 23 to 26 combine the three mechanisms previously used. Table 6 details these combinations and the configurations of CPU, nice-based prioritization, and bandwidth control. In all cases, the priority slice (ice 1) was configured with the prioritization used in Experiment 10, that is, with nice set to -5, while the other slices were set to a value of 5. The CPU limit follows the configurations from Experiments 6 and 7, with a CPU limit of 1000 milicpu for all priority slices. CPU limits of 500 milicpu were applied in Experiments 25 and 26. Regarding bandwidth limitation, in Experiments 24 and 26, a bandwidth limit of 150 Mbps was assigned, while in Experiments 25 and 26, a CPU limit of 500 milicpu was defined for the non-priority slices.

### v2

@@ 1 Introduction
In this section, we review the main changes introduced by 5G networks, focusing on the network core, including the network functions for both the control plane and the user plane, as well as the isolation mechanisms and innovations that differentiate it from previous generations. We also present the main open-source 5G applications. The goal is to understand the operation of 5G, with an emphasis on network isolation, providing the foundation for building an adequate 5G network for experimentation. Finally, we present related work on the topic of 5G network slice isolation.
@@ 2.6.1 UERANSIM
UERANSIM is a simulator for user equipment (UE) and radio access network (RAN) in 5G networks, designed as an open-source solution that simulates both 5G mobile devices and base stations. The project aims to enable testing and research on 5G networks. However, UERANSIM does not offer a complete implementation of the 5G physical layer, it have limitations in radio channel and physical layer modeling [18]. The 5G-NR radio interface is only partially implemented and simulated using the UDP protocol [19]. The simulator includes the components nr-gnb and nr-ue, which represent, respectively, the gNodeB (gNB) and the UE. When establishing a connection with the 5G network, the UE creates a network interface responsible for tunneling with the user plane function (UPF).
@@ 2.7 Tool Comparisons
3 Related Work
Several studies have explored 5G slicing testbeds and mechanisms to ensure performance and isolation guarantees. Esmaeily and Kralevska [21] provide a comprehensive comparison of 5G slicing environments, defining primary criteria such as support for SDN, NFV, and cloud computing, as well as dynamic orchestration and multi-tenancy. Secondary criteria include multi-RAT support, E2E slicing, ML integration, and open-source availability. Despite these advances, the authors observe that most evaluated testbeds offer only partial isolation or limit slicing to specific protocol layers.
Rouili et al. [18] evaluate the fidelity of open-source 5G standalone testbeds by comparing simulation/emulation tools with SDR-based deployments. Although their focus is on RAN performance rather than core-level slicing, they underscore the challenges of achieving realistic 5G environments using open-source components—challenges that similarly affect core slicing deployments. Similarly, Garcia-Aviles et al. [22, 23] propose the POSENS protocol, which enables E2E slicing through modifications to PDCP and RRC layers and traffic redirection at a load balancer node to enforce bandwidth limitations between slices. In contrast, Shorov et al. [9] and Dzogovic et al. [24] present early testbeds focused on slice selection and deployment across virtualized cores, though without addressing slice isolation explicitly.
More recent work investigates slicing from a performance assurance perspective. Nikolaidis et al. [25] address per-user QoS guarantees in RAN slices using a multiclass queuing model for bandwidth provisioning. While their focus is on intra-slice differentiation and SLA negotiation, our work targets inter-slice performance isolation using infrastructure-level controls. De Simone et al. [26] present a formal approach to optimizing performance and availability in resilient 5G architectures through stochastic modeling and empirical validation using Open5GS. Unlike their SLA-driven redundancy planning, we concentrate on mitigating slice interference through direct resource controls in edge/cloud-native environments.
Other studies adopt broader system-level perspectives. Minardi et al. [27] explore SDN-based control strategies for integrated terrestrial and non-terrestrial networks, focusing on reactive and proactive traffic rerouting rather than resource isolation. Adeppady et al. [28] propose the iPlace heuristic to reduce microservice interference and deployment time in constrained cloud environments. Their work highlights the importance of resource-aware orchestration, aligning with our focus on runtime performance stability in private 5G infrastructures. Muro et al. [29] investigate the impact of the Noisy Neighbour phenomenon in virtualized 5G cores and apply machine learning for detection and forecasting. While they propose reactive mitigation, our work emphasizes proactive enforcement of CPU and bandwidth limits to avoid contention altogether.
At the orchestration layer, Ait Aba et al. [30] enhance Kubernetes deployment strategies by integrating Genetic Algorithms to solve the Virtual Network Embedding problem and reduce failed slice deployments. Whereas their contribution targets slice admission efficiency, our approach deals with maintaining performance guarantees post-deployment, reinforcing the need to adapt general-purpose orchestrators to telecom-specific requirements.
Taken together, these studies demonstrate the growing interest in network slicing and resource management for 5G. However, they also reveal gaps in effective, reproducible solutions for slice isolation in open-source, resource-constrained environments. Our work contributes to this space by experimentally evaluating native resource control mechanisms—such as CPU quotas, bandwidth shaping, and prioritization—in Kubernetes-managed private 5G deployments, offering a practical and reproducible path toward performance-aware slice management.
@@ 5.1 Experimentation Methodology
During the experiments, several UEs (or pods) are executed over a 20-minute period. Initially, to represent the high-priority slice (Slice 1), 4 pods are started simulating 4 cameras streaming high-quality video at an operating room. Video is streamed over the edge-located UPF and delivered to a simulated cloud service (iPerf server). UEs connected to lower-priority slices (Slices 2 to 5) serve as background traffic and generate the same traffic patterns as the priority ones. In Slices 2 to 5, at every 4 minutes of experimentation, a new UE is added, while the pods are allocated to the node corresponding to the slice, as illustrated in Figure 3. Video transmission is simulated using the isochronous [^6] https://manpages.debian.org/bullseye/iperf/iperf.1.en.html option of IPerf, configured with 60 frames per second at a bitrate of 45 Mbits per second, and no variation in frame size.
@@ 5.8 Experiments with CPU limitation, bandwidth limitation, and prioritization
Experiments 23 to 26 combine the three mechanisms previously used. Table 6 details these combinations and the configurations of CPU, nice-based prioritization, and bandwidth control. In all cases, the priority slice (Slice 1) was configured with the prioritization used in Experiment 10, that is, with nice set to -5, while the other slices were set to a value of 5. The CPU limit follows the configurations from Experiments 6 and 7, with a CPU limit of 1000 milicpu for all priority slices. CPU limits of 500 milicpu were applied in Experiments 25 and 26. Regarding bandwidth limitation, in Experiments 24 and 26, a bandwidth limit of 150 Mbps was assigned, while in Experiments 25 and 26, a CPU limit of 500 milicpu was defined for the non-priority slices.